# Scattering Theory of Chiral Edge Modes in Topological Magnon Insulators

### v1

@@ Abstract
Topological magnon insulators exhibit robust edge modes with chiral properties similar to quantum Hall edge states. However, due to their strong localization at the edges, interactions between these chiral edge magnons can be significant, as we show in a model of coupled magnon-conserving spin chains in an electric field gradient. The chiral edge modes remain edge-localized and do not scatter into the bulk, and we characterize their scattering phase: for strongly-localized edge modes we observe significant deviation from the bare scattering phase. This renormalization of edge scattering can be attributed to bound bulk modes resonating with the chiral edge magnons, in the spirit of Feshbach resonances in atomic physics. We argue a real-time measurement protocol using spin-polarized scanning tunneling spectroscopy to study their scattering dynamics. Our result show that interaction among magnons can be encoded in an effective edge model of reduced dimensionality, where the interactions with the bulk renormalize the effective couplings. This work paves the way to develop a many-body effective theory for chiral edge magnons.
@@
Introduction – Quantum magnets with a topologically trivial ground state supporting topologically nontrivial spin excitations are an emerging platform to study the topology of collective excitations [1]. They are found in ordered magnets [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and quantum paramagnets [15, 16, 17, 18, 19]. When time-reversal symmetry is broken, chiral spin excitations propagating along the boundaries of the sample are protected by the topology of the bulk, giving rise to the idea of topological magnonics [20, 21, 22, 23, 24]. Although direct experimental evidence for chiral edge spin excitations is yet to be achieved, transport [3, 25, 26, 27] and neutron scattering data [28, 29, 30, 31, 32] compatible with topological spin excitations have been reported. However, the stability against decay of topological magnon modes still needs to be firmly established [33, 34, 35]. The properties of magnon edge modes can be investigated by spin-polarized scanning tunneling spectroscopy (STS) through local excitation [36, 37]. While much of the focus thus far has been on investigating the properties of isolated edge excitations, many experimental setups operate in a regime where multiple edge modes are excited simultaneously. As they will interact with one another, it is an important open challenge to study the characteristics of scattering between edge modes. A positive answer would lay the groundwork for building an effective edge theory at the many-body level.
Herein, we show the stability of chiral magnons upon scattering and characterize their properties. We emphasize the critical role of the bulk spectrum in determining the interaction characteristics of edge states. Scattering of chiral magnons can be directly explored using spin-polarized STS through a real-time injection measurement protocol, as shown in Fig. 1 a): Chiral edge magnons with different velocities can be sequentially excited. As illustrated in Fig. 1 b), the interactions between edge modes are reflected in a scattering delay $\Delta d$, measured relative to the evolution of free particles. This delay can be extracted by monitoring the chiral excitations at a different position of the edge, see Fig. 1 a).
@@
Figure 1: Topological magnons in coupled wires. a) Sketch of the model (1): we consider a two-dimensional system made of an array of spin1/212-1/2- 1 / 2 chains. The presence of a linear Aharanov-Casher (AC) phase induced by a constant electric field gradient in the $y$ direction $\vec{E}=\Delta E\vec{y}$, gives rise to localized bulk modes and chiral edge modes. b) Cartoon scattering of two chiral magnons at the same edge. Our results demonstrate that the scattering of edge modes is elastic (no scattering into the bulk), but resonances with bulk modes strongly renormalize effective interactions, manifesting in a scattering shift $\Delta d$.
@@
The model – We require both spin conservation and ferromagnetism to ensure the number of magnons is constant. We assume that any perturbation deviating from this condition—setting a finite timescale for the proposed scattering experiments—is small enough to be negligible. As a concrete example, we construct a two-dimensional system from exactly solvable one-dimensional XXZ spin chains with tunable magnon-magnon interactions [38, 39, 40]. The individual chains are arranged in the $x$-$y$ plane, aligned parallel in the $x$-direction, and labeled according to their equidistant positions along the $y$-axis, see Fig. 1 a). The Hamiltonian for the $y$-th chain takes the form
$\mathcal{H}^{(y)}=-\frac{J_{x}}{2}\sum_{x}\Bigg{[}\Big{(}e^{-i\varphi_{y}}\sigma_{x,y}^{+}\sigma_{x+1,y}^{-}+\mathrm{h.c.}\Big{)}+\Delta\sigma_{x,y}^{z}\sigma_{x+1,y}^{z}\Bigg{]}.$ (1)
@@
The interactions between magnons are controlled by the anisotropy $\Delta$, with magnons behaving as hardcore bosons in the limit $\Delta=0$. To induce chiral edge modes, we use the Aharanov-Casher (AC) phase [41, 42, 43, 44, 45, 46]: an electric field gradient causes a $y$-dependent phase $\varphi_{y}\propto\int_{(x,y)}^{(x+1,y)}\mathrm{d}\vec{x}\cdot(\vec{E}\wedge\hat{e}_{z})\equiv y\delta\varphi$ for the intra-chain hopping. We measure energy scales in units of the intra-wire hopping and hence set $J_{x}=1$. The Hilbert space for Eq. (1) contains sectors with a definite number of magnons, and the single magnon spectrum is obtained by Fourier transformation. In the absence of any interchain coupling, the phase $\varphi_{y}$ is equivalent to different gauges of lattice momenta, thus resulting in a global shift of the one-magnon bands $k\to k+\varphi_{y}$, see Fig. 2 a) (gray lines).
@@
Figure 2: Real time scattering of chiral magnons. a) Single magnon spectrum with $\delta\varphi=2\pi/5$ for the case of uncoupled wires, consisting of multiple copies of single-wire bands shifted in the momentum space (gray lines). Inter-wire hopping $J_{y}\neq 0$ gives band repulsion, and the spectrum shifts to stationary bulk modes and edge chiral states. b) Real-time simulation of a scattering event of two chiral magnons with total momentum $K=-0.39$ and relative momentum $q=-0.20$, and parameters $(J_{y},\Delta)=(0.5,0.8)$. We track the bulk-integrated magnon distribution $|\psi(x)|^{2}$, with $x$ the relative distance. For a better visualization, we consider a comoving frame with the same velocity of the pair of magnons. The gray dashed lines are the trajectories of the initial position of the wavepackets, evolved with the initial velocities (free evolution). c) The state remains localized at the edge during the scattering, as shown in the average position in the transverse direction $\left\langle y\right\rangle$. d) We compare the free evolution (gray dashed line) with the peaks position (blue line), showing displacement after scattering.
We now allow for tunneling between the chains leading to the full Hamiltonian
$\mathcal{H}=\sum_{y}\mathcal{H}^{(y)}-J_{y}\sum_{x,y}\frac{1}{2}\Big{(}\sigma_{x,y}^{+}\sigma_{x,y+1}^{-}+\mathrm{h.c.}\Big{)}.$ (2)
Figure 3: Scattering phases. a) Left: Allowed energy-momentum window for scattering magnons (blue) embedded in the two magnon continuum (gray area). Non-dispersive bulk bound modes with neighboring magnons (BS) are partially overlapping with the energy window of the asymptotic edge magnons (blue). The parameters are $(J_{y},\Delta)=(0.5,0.0)$. Right: Momentum-integrated density of states capturing the low density of states within the bulk gap (lightblue shaded region). b) Extracted scattering phases for asymptotic states of different relative momenta $q$ for fixed total momenta $K\in\{-0.43,0,0.43\}$ (left, middle, right column) for different $\Delta\in\{0.1,0.3,1.0\}$. The one-dimensional XXZ prediction (gray line) is provided for comparison, highlighting a strong renormalization for positive $K$ and small $\Delta$. c) Comparison of effective interaction strength $c$ extracted from the scattering phase and chiral dispersion for numerically extracted results and analytical predictions for one-dimensional XXZ result (gray lines).
@@
A nonzero inter-wire hopping $J_{y}$ causes level repulsion between the wires’ energy bands, opening up a bulk gap in the system. Only the branches associated with the edge wires escape level repulsion in certain momenta regions $k\in[k_{1},k_{2}]$, giving rise to localized chiral edge modes with dispersion $\varepsilon(k)$, and group velocity $v(k)\equiv\partial_{k}\varepsilon(k)$ [47]. In the thermodynamic limit, bulk modes become localized and have flat energy bands, while edge modes (i) have a well-defined chirality given by the sign of the group velocity, and (ii) are dispersive, i.e. $\partial_{k}v(k)\neq 0$; thus they scatter.
@@
We show an example of the chiral spectrum in Fig. 2 a) for inter-wire hopping strength $J_{y}=0.5$, and indicate the localization of individual magnon modes in the bulk by color coding according to Fig. 1 a), for $\delta\varphi=\frac{2\pi}{5}$. For the sake of consistency, we fix the AC phase to this value throughout the rest of this work, but generalization to other values of $\delta\varphi$ are straightforward and lead to similar results.
Real-time scattering of Chiral Magnons – When scattering channels into bulk modes or other edge bands are closed, the scattering remains effectively one-dimensional. As a result, due to energy and momentum conservation, two-magnon scattering is elastic, preserving the individual momenta. However, this does not translate into a trivial scattering event as interactions induce a scattering shift $\Delta d$, which represents the spatial delay of the interacting particles compared to the free propagation of wave packets. This delay provides a direct measure of the strength of interactions between excitations. We now investigate real-time scattering of edge modes using exact diagonalization in the two magnon subspace. The bulk-integrated magnon distribution $|\psi(x)|^{2}$ as a function of time is shown in Fig. 2 b). To confirm the absence of leakage into bulk modes or transitions to other edge states during scattering, we compute the wave packets’ average edge localization $\left\langle y\right\rangle$, which stays constant in time, see Fig. 2 c). In Fig. 2 d), we quantify the scattering shift $\Delta d$ by comparing the actual peak positions (solid blue line) with those predicted for free propagation (dashed gray line). While this real-time scattering analysis provides a way to characterize the two-body problem at the edge, we will next develop a more efficient method to extract the scattering information encoded in $\Delta d$ across a broader parameter range. This approach will be contrasted against predictions from the standard XXZ theory on the edge wires.
Characterization of edge scattering phases – We directly extract the desired scattering information from the interacting two particle spectrum. First, we utilize translational invariance along the chain direction, allowing us to represent (2) in momentum space. We then focus on the two-magnon sector of the Hilbert space and diagonalize the Hamiltonian within each center-of-mass (COM) momentum sector $K$. An example of the momentum-resolved low-energy continuum spectrum is shown in Fig. 3 a) (gray shaded region) for parameters $J_{y}=0.5,\Delta=0$. The topological bulk gap is visible in the momentum-integrated density of states (right panel). For our scattering analysis, we concentrate on the energy-momentum region allowed for two chiral magnons localized at the lower edge, as highlighted by the blue shaded area in Fig. 3 a). The energy-momentum window $[E_{1}(K),E_{2}(K)]$ for such an asymptotic state of two edge magnons with momenta $k_{1}$ and $k_{2}$ is determined from the chiral magnon dispersion, i.e. $E_{1}(K)\leq\varepsilon(k_{1})+\varepsilon(k_{2})\leq E_{2}(K)$, with $k_{1}+k_{2}=K$. The relevant asymptotic states for total momentum $K$ are generally described by a superposition of plane waves $\psi_{K}(x,y_{1},y_{2})\propto e^{i(k_{1}-k_{2})x/2}\bar{\psi}_{k_{1}}(y_{1})\bar{\psi}_{k_{2}}(y_{2})+e^{-i(k_{1}-k_{2})x/2}e^{i\phi(k_{1},k_{2})}\bar{\psi}_{k_{1}}(y_{2})\bar{\psi}_{k_{2}}(y_{1})$ representing the real-space wave function for magnons with relative distance $x$ in wires $y_{1}$ and $y_{2}$. In this representation, we ignore the trivial global phase contribution from the COM motion with momentum $K$. Fitting the actual wavefunction for a given total momentum $K$ to this functional form allows us to extract the scattering phase $\phi$ as a function of the relative momentum $q=(k_{1}-k_{2})/2$ (see Supplementary Material for details). The scattering phase $\phi$ connects back to the real-time scattering results in Fig. 2 d) through its relation to the scattering shift, i.e. $\Delta d\propto\partial_{k_{1}}\phi(k_{1},k_{2})$ [48].

### v2

@@ Abstract
Topological magnon insulators exhibit robust edge modes with chiral properties similar to quantum Hall edge states. However, due to their strong localization at the edges, interactions between these chiral edge magnons can be significant, as we show in a model of coupled magnon-conserving spin chains in an electric field gradient. The chiral edge modes remain edge-localized and do not scatter into the bulk, and we characterize their scattering phase: for strongly-localized edge modes we observe significant deviation from the bare scattering phase. This renormalization of edge scattering can be attributed to bound bulk modes resonating with the chiral edge magnons, in the spirit of Feshbach resonances in atomic physics. We argue that the scattering dynamics can be probed experimentally with a real-time measurement protocol using inelastic scanning tunneling spectroscopy. Our results show that interaction among magnons can be encoded in an effective edge model of reduced dimensionality, where the interactions with the bulk renormalize the effective couplings. Our work introduces a systematic way to determine the many-body effective theory for edge states in topological magnon insulators.
I Introduction
@@
Quantum magnets with a topologically trivial ground state supporting topologically nontrivial spin excitations are an emerging platform to study the topology of collective excitations [1]. They are found in ordered magnets [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and quantum paramagnets [15, 16, 17, 18, 19]. When time-reversal symmetry is broken, chiral spin excitations propagating along the boundaries of the sample give rise to the idea of topological magnonics [20, 21, 22, 23, 24]. Although direct experimental evidence for chiral edge spin excitations is yet to be achieved, transport [3, 25, 26, 27] and neutron scattering data [28, 29, 30, 31, 32] compatible with topological spin excitations have been reported. However, the stability against decay of topological magnon modes still needs to be firmly established [33, 34, 35]. The properties of magnon edge modes can be investigated by inelastic scanning tunneling spectroscopy (STS) through local excitation [36, 37]. While much of the focus thus far has been on investigating the properties of isolated edge excitations, many experimental setups operate in a regime where multiple edge modes are excited simultaneously. For example, spintronics or general transport experiments investigate spin and energy transport governed by an ensemble of edge modes with various energies and velocities. Naturally, they will interact with one another and consequences of interactions are expected to be reflected in transport properties of the system. It is therefore an important open challenge to study the characteristics of interactions between edge modes.
@@
Figure 1: Topological magnons in coupled wires. a) Sketch of the model (2): we consider a two-dimensional system made of an array of spin1/2-1/2 chains. The presence of a linear Aharonov-Casher (AC) phase induced by a constant electric field gradient in the $y$-direction $\vec{E}=\Delta E\vec{y}$, gives rise to localized bulk modes and chiral edge modes. b) Cartoon scattering of two chiral magnons at the same edge. Our results demonstrate that the scattering of edge modes is elastic (no scattering into the bulk), but resonances with bulk modes strongly renormalize effective interactions, manifesting in scattering shifts $d_{1},d_{2}$ of the individual edge modes.
In this work, we consider a two-dimensional lattice of spins in which edge magnons are induced by the Aharonov-Casher (AC) effect [38, 39, 40, 41, 42, 43]; see Fig. 1 (a). We use this tunable toy model as a proof of principle to develop a protocol for constructing an effective interacting edge Hamiltonian, thereby highlighting the peculiar role of resonances with bulk excitations for renormalization of interactions at the edge. We show that scattering of edge magnons remains confined to the edge with no leakage into bulk degrees of freedom, and we capture their dynamics within a two-body effective edge Hamiltonian
where $k_{1,2}$ are the edge-momenta of the magnons, and $\varepsilon(k)$ is the single-magnon energy. The effective interaction $U_{\text{2-eff}}(k_{1},k_{2})$ is determined by the scattering phase $\phi(k_{1},k_{2})$ [44, 45, 46], which encodes the additional phase a two-magnon wavefunction acquires from scattering. Interestingly, we find $\phi(k_{1},k_{2})$ to be noticeably renormalized by virtual processes with bi-magnon bulk bound states.
Our work paves the way for a systematic development of an interacting theory of chiral edge magnons, and it provides the groundwork to use powerful theoretical methods of one-dimensional quantum systems to deepen our understanding of two-dimensional topological magnets. Our results are pertinent for experimental studies: we discuss how scattering of chiral magnons can be explored using inelastic STS through a real-time injection measurement protocol, as shown in Fig. 1 (a), where the presence of a non-trivial scattering phase manifests as a detectable shift of excitations after scattering; see Fig. 1 (b). Furthermore, understanding edge interactions could prove instrumental for future coherent magnon-based quantum technologies, such as magnonic interferometers and quantum circulators.
II The model
@@
We consider a ferromagnet with spin-conservation, ensuring magnons are stable excitations, and assume that any deviation from this condition can be neglected on the time scale of scattering experiments. As a toy model, we build a two-dimensional system from exactly solvable one-dimensional XXZ spin chains with tunable magnon-magnon interactions [47, 48, 49]. The individual chains are arranged in the $x$-$y$ plane, aligned parallel in the $x$-direction, and labeled according to their equidistant positions along the $y$-axis, see Fig. 1 (a). The Hamiltonian for the $y$-th chain is
$\mathcal{H}^{(y)}=-\frac{J_{x}}{2}\sum_{x}\big(e^{-i\varphi_{y}}\sigma_{x,y}^{+}\sigma_{x+1,y}^{-}+\mathrm{h.c.}\big)+\Delta\sigma_{x,y}^{z}\sigma_{x+1,y}^{z}.$ (2)
@@
The magnons’ interactions are controlled by the anisotropy $\Delta$, with magnons behaving as hardcore bosons in the limit $\Delta=0$. Next we stabilize a magnetically ordered state along the $z$-direction for all $\Delta$ via an external magnetic field. Chiral edge magnons can then be induced on top of this ordered state via an Aharonov-Casher (AC) phase [38, 39, 40, 41, 42, 43]: an electric field gradient along the $y$-direction causes a chain-dependent phase $\varphi_{y}\propto\int_{(x,y)}^{(x+1,y)}\mathrm{d}\vec{x}\cdot(\vec{E}\wedge\hat{e}_{z})\equiv y\delta\varphi$ for the intra-chain hopping. We measure energy scales in units of the intra-wire hopping and set $J_{x}\!=\!1$. The eigenstates of Eq. (2) have a definite number of magnons. In the absence of interchain coupling, the phase $\varphi_{y}$ is equivalent to different gauges of lattice momenta, resulting in a global shift of the one-magnon bands $k\to k+\varphi_{y}$, see Fig. 2 (a) (gray lines).
@@
Figure 2: Real time scattering of chiral magnons. a) Single magnon spectrum with $\delta\varphi=2\pi/5$ for the case of uncoupled wires, consisting of multiple copies of single-wire bands shifted in the momentum space (gray lines). Inter-wire hopping $J_{y}\neq 0$ causes band repulsion, and the spectrum shifts to stationary bulk modes and chiral edge bands (within gray box). Location of the individual bands is indicated by color coding. b) Real-time simulation of a scattering event of two chiral magnons with total momentum $K=-\frac{\pi}{8}$ and relative momentum $q=-\frac{\pi}{16}$, and parameters $(J_{y},\Delta)=(0.5,0.4)$. We track the bulk-integrated magnon distribution $|\psi(x)|^{2}$, with $x$ the relative distance. For a better visualization, we consider a comoving frame with the average velocity of the pair of magnons. The gray dashed lines are the trajectories of the initial position of the wavepackets, evolved with the initial velocities (free evolution). c) The state remains localized at the edge during the scattering, as shown in the average position in the transverse direction $\left\langle y\right\rangle$, where $\left\langle y\right\rangle\!=\!1$ corresponds to magnons being located in the lowest chain. d) We compare the free evolution (gray dashed line) with the peak positions (blue line), showing displacements $d_{1},d_{2}$ after scattering.
In the full Hamiltonian,
$\mathcal{H}=\sum_{y}\mathcal{H}^{(y)}-J_{y}\sum_{x,y}\frac{1}{2}\Big(\sigma_{x,y}^{+}\sigma_{x,y+1}^{-}+\mathrm{h.c.}\Big),$ (3)
@@
the inter-wire hopping $J_{y}$ causes level repulsion between the wires’ energy bands, opening up a bulk gap in the system. Only the energy branches associated with the edge wires escape level repulsion in certain momenta regions $k\in[k_{1},k_{2}]$, giving rise to localized chiral edge modes with dispersion $\varepsilon(k)$, and group velocity $v(k)\!\equiv\!\partial_{k}\varepsilon(k)$ [50]. In the thermodynamic limit, bulk modes become localized and have flat energy bands, while edge modes have a well-defined chirality given by the sign of the group velocity, and they are dispersive, i.e. $\partial_{k}v(k)\!\neq\!0$, thus they scatter. We emphasize that also Ising inter-wire interactions can be added to Eq. (3). While this leads to no qualitative changes in the physics it complicates the following discussion of bulk resonances and is hence only examined in Appendix E.
@@
We show an example of the chiral spectrum in Fig. 2 (a) for inter-wire hopping strength $J_{y}\!=\!0.5$, and we indicate the location of individual magnon modes in the bulk by color coding according to the sketch of Fig. 1 (a), for $\delta\varphi=\frac{2\pi}{5}$. The chiral dispersion $\varepsilon(k)$ is located within the gray box. We fix the AC phase to this value throughout the rest of this work, but other values are straightforward and lead to similar results.
III Real-Time Scattering of Chiral Magnons
When scattering channels into bulk modes or other edge bands are closed, the scattering remains effectively one-dimensional. Due to energy and momentum conservation, two-magnon scattering is elastic preserving the individual momenta, but it remains highly non-trivial as interactions induce scattering shifts $d_{1},d_{2}$, displacing the magnons from their free trajectories. This delay quantifies the strength of interactions between excitations. We investigate this effect in real-time scattering of edge modes using exact diagonalization in the two magnon subspace. For this, we implement an initial state of two localized magnons of given total momentum $K$ and relative momentum $q$ separated from each other in the lattice. To get optimal overlap of our initial state with the eigenstates of the system, we construct the real space wavefunction from the single magnon eigenstates $\bar{\psi}_{k_{1}}(y_{1}),\bar{\psi}_{k_{2}}(y_{2})$ of the Hamiltonian with momenta $k_{1}=K/2+q$ and $k_{2}=K/2-q$. In real space, single magnon eigenstates are given by plane waves of momentum $k_{1}$ and amplitude $\bar{\psi}_{k_{1}}(y_{1})$ in wire $y_{1}$. We localize both single magnon states by sufficient dressing with Gaussian filters. The overall wavefunction $\psi_{k_{1},k_{2}}(x_{1},x_{2},y_{1},y_{2})$ is hence given as a superposition of two single magnon states, i.e.
$\psi_{k_{1},k_{2}}(x_{1},x_{2},y_{1},y_{2})\propto\;$ $\bar{\psi}_{k_{1}}(y_{1})\bar{\psi}_{k_{2}}(y_{2})G_{\sigma X_{1}}(x_{1})G_{\sigma X_{2}}(x_{2})$
$\times e^{-i(k_{1}x_{1}+k_{2}x_{2})}$
$\text{with}\quad G_{\sigma X_{0}}(x)$ $\equiv\dfrac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{1}{2\sigma^{2}}(x-{X}_{0})^{2}}.$ (4)
Scattering between the individual magnon wavepackets initially localized at positions $X_{1},X_{2}$ with width $\sigma$ is then studied using standard time evolution under the Hamiltonian of Eq. (3). The results of an exemplary evolution are summarized in Fig. 2 (b) - (d). The bulk-integrated magnon distribution $|\psi(x)|^{2}$ as a function of time is shown in Fig. 2 (b). To confirm the absence of leakage into bulk modes or transitions to other edge states during scattering, we compute the wave packets’ average edge localization $\left\langle y\right\rangle$, which stays constant in time, see Fig. 2 (c). In Fig. 2 (d), we quantify the scattering shift $d_{1},d_{2}$ by comparing the actual peak positions (solid blue line) with those predicted for free propagation (dashed gray line). For details on the numerical implementation see Appendix B. While the scattering shift observed in real-time evolution is a direct manifestation of interactions, a more systematic study is provided by the scattering phase $\phi(k_{1},k_{2})$ which is directly connected with the scattering shift as $d_{1,2}=\partial_{k_{1,2}}\phi(k_{1},k_{2})$ [51].
Figure 3: Scattering phases. a) Left: Allowed energy-momentum window for scattering magnons at the lower edge (blue) embedded in the two-magnon continuum (gray area). The asymptotic energy region (blue) overlaps with non-dispersive bulk bands associated to bound states of two neighboring magnons (orange). The parameters are $(J_{y},\Delta)\!=\!(0.5,0.0)$. Right: Momentum-integrated two-magnon density of states capturing the low density of states within the bulk gap (lightblue shaded region). b) Extracted scattering phases for asymptotic states of different relative momenta $q$ for fixed total momenta $K\in\{-0.43,0,0.43\}$ (left, middle, right column) for selected $\Delta\in\{0.1,0.3,1.0\}$. The one-dimensional XXZ prediction (gray line) is provided for comparison, highlighting a strong renormalization for positive $K$ and small $\Delta$. c) Comparison of equal-momenta effective interaction strength $U_{\text{2-eff}}(K/2,K/2)$ extracted from the scattering phase and chiral dispersion $\varepsilon(k)$ for numerically extracted results and analytical predictions for one-dimensional XXZ (gray lines).
IV Characterization of Edge Scattering Phases
When two wavepackets of edge magnons collide, they accumulate a scattering phase $\phi(k_{1},k_{2})$ depending on the asymptotic momenta $k_{1,2}$. For non-interacting particles, the scattering phase $\phi(k_{1},k_{2})$ vanishes; for hard-core bosons, $\phi(k_{1},k_{2})$ approaches $\pi$; and for genuinely interacting cases, it exhibits a non-trivial momentum dependence. The scattering phase determines the effective interactions $U_{\text{2-eff}}(k_{1},k_{2})$ in Eq. (1) through the inverse scattering method [44, 45, 46] : the solution is unique, and can be numerically found evaluating the Gel’fand-Levitan-Marchenko’s equations [45] or solving an optimization problem based on chosen trial potentials [46]. Instead of aiming for a full characterization of the effective interactions we can, however, focus on the emergent theory in the assumption that the momenta of the scattering magnons belong to a small shell $k\in I_{p}\equiv[p-\delta p,p+\delta p]$. In this approximation, we Taylor expand the dispersion law
$\varepsilon(k_{1,2})\simeq\varepsilon(p)+v(p)(k_{1,2}-p)+\frac{1}{2m^{*}(p)}(k_{1,2}-p)^{2}+\cdots,$ (5)
where we introduced the velocity $v(k)=\partial_{k}\varepsilon(k)$ and effective mass $m^{*}(k)\equiv[\partial_{k}^{2}\varepsilon(k)]^{-1}$, and terms of order $\mathcal{O}\big((k_{1,2}-p)^{3}\big)$ are not shown. Next, we approximate the interaction $U_{2-\text{eff}}(k_{1},k_{2})\simeq U_{2-\text{eff}}(p,p)$. Within this approximation, it is easy to convert $H_{\text{2-eff}}$ into a real space Hamiltonian, where the spatial coordinates $x_{1,2}$ conjugated to the small momenta shift $k_{1,2}-p$ are approximated as continuous
$\Delta\mathcal{H}_{\text{2-eff}}\simeq$ $-\tfrac{1}{2m^{*}(p)}(\partial^{2}_{x_{1}}+\partial^{2}_{x_{2}})+iv(p)(\partial_{x_{1}}+\partial_{x_{2}})$
$+c(p)\delta(x_{1}-x_{2})\,$, (6)
where $\Delta\mathcal{H}_{\text{2-eff}}=\mathcal{H}_{\text{2-eff}}-2\varepsilon(p)$ and $c(p)\equiv U_{\text{2-eff}}(p,p)$. The Hamiltonian (IV) describes the two-particle sector of the Lieb-Liniger (LL) model [52, 53]: with a contact potential, in a moving frame with velocity $v(p)$. The scattering phase of this model is [52, 53]
$\phi_{\mathrm{LL}}(k_{1},k_{2})=-2\arctan\left(\tfrac{k_{1}-k_{2}}{m^{*}(p)c(p)}\right).$ (7)
By comparing the true scattering phase with the effective one, we can determine the effective interaction. Since we are within a small relative momentum approximation, we fix the effective interaction $c(p)$ through the scattering length $a(p)\equiv\lim_{k_{1}\to p}\partial_{k_{1}}\phi(k_{1},p)$, and the LL correspondence $a(p)=-1/(m^{*}(p)c(p))$. This, in particular, implies $U_{\text{2-eff}}(k,k)=-1/[m^{*}(k)a(k)]$.
Quantifying the effective theory at the edge hence requires characterization of the scattering phase $\phi(k_{1},k_{2})$. As an alternative to the real-space scattering protocol presented above, we systematically probe the scattering phase directly through the eigenstates of the two-dimensional problem, by searching the spectrum for scattering states of chiral edge modes. In the asymptotic region $x_{1}\ll x_{2}$, with $x_{1,2}$ the position of the magnons in the edge direction, the eigenfunction reads
$\psi(x_{1},x_{2},y_{1},y_{2})$ $\propto e^{ix_{1}k_{1}+ix_{2}k_{2}}\bar{\psi}_{k_{1}}(y_{1})\bar{\psi}_{k_{2}}(y_{2})$
$+e^{i\phi(k_{1},k_{2})}e^{ix_{1}k_{2}+ix_{2}k_{1}}\bar{\psi}_{k_{1}}(y_{2})\bar{\psi}_{k_{2}}(y_{1}),$ (8)
where $e^{ikx}\bar{\psi}_{k}(y)$ is the one-magnon wavefunction.